# Searching in Euclidean Spaces with Predictions

### v1

@@ Abstract
We study the problem of searching for a target at some unknown location in $\mathbb{R}^{d}$ when additional information regarding the position of the target is available in the form of predictions. In our setting, predictions come as approximate distances to the target: for each point $p\in\mathbb{R}^{d}$ that the searcher visits, we obtain a value $\lambda(p)$ such that $|p\bm{t}|\leq\lambda(p)\leq c\cdot|p\bm{t}|$, where $c\geq 1$ is a fixed constant, $\bm{t}$ is the position of the target, and $|p\bm{t}|$ is the Euclidean distance of $p$ to $\bm{t}$. The cost of the search is the length of the path followed by the searcher. Our main positive result is a strategy that achieves $(12c)^{d+1}$-competitive ratio, even when the constant $c$ is unknown. We also give a lower bound of roughly $(c/16)^{d-1}$ on the competitive ratio of any search strategy in $\mathbb{R}^{d}$.
@@ 1 Introduction
The problem of searching for a target positioned at some unknown location in some region is a classic search game problem that has been well-studied in both fields of Computational Geometry and Operations Research. The problem comes in many different versions including, linear search (i.e., searching for a target on a line) [9, 6, 7, 15], searching in the plane [8, 20, 16, 4], searching in concurrent rays [14, 15, 4], and searching inside polygonal regions [22, 28]. The book by Alpern and Gal [1] provides an extensive overview of general search games, while Ghosh and Klein [17] survey search problems in planar domains.
@@ 1.1 Problem setup
As we search for the target, we have approximate information about the distance to it from each point that we have visited in so far. More precisely, we assume that there is a constant $c\geq 1$ and an unknown function
@@ 1.2 Our contribution
  • We show that for each dimension $d$ and each constant prediction factor $c\geq 1$ there is a search strategy with competitive ratio $(6c)^{d+1}/3$. To achieve this, we use $\varepsilon$-nets from metric spaces, also known as $r$-nets, and provide a path of finite length but an infinite number of pieces. This result holds assuming that we know the prediction factor $c$. For unknown prediction factor, a similar statement holds with a slightly different search strategy and larger competitive ratio. This result is given in Section 4.
  • We show that for $c>4$, any deterministic search strategy in $\mathbb{R}^{d}$ with $c$-predictions will have a competitive ratio of at least $(1/4)\cdot(c/16)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$. For this result we again use $\varepsilon$-nets as a basic tool and we employ techniques developed for the Euclidean Traveling Salesperson with Neighbourhoods. With a slightly worse constant, the lower bound holds also for randomized search strategies. This lower bound is shown in Section 5.
@@
To our knowledge, this is the first search problem in Euclidean spaces with $d\geq 2$ where a search point reaches a target point with constant competitive ratio; see the discussion below about related work. For our upper and lower bounds we use bounds on the cardinality of $\varepsilon$-nets. For the sake of simplicity in the calculations, we used suboptimal but simple-to-parse estimates in the simplifications and the cardinality of $\varepsilon$-nets. In any case, our upper and lower bounds for the best competitive ratio are still far apart.
@@ 1.3 Related work
In another variant of the problem in the plane by Gal [16], the searcher travels along a path until the target lies on the segment connecting the searcher’s current and starting positions, essentially sweeping around its starting position with an infinitely elastic cord until the target is swept. For this problem, Gal [16] gave a spiral search strategy achieving a competitive ratio of $17.289$ while Langetepe [24] showed that this ratio is optimal.
@@ 2 Notation and preliminaries
The following bound on $\varepsilon$-nets follows from a well-known technique using volumes.
@@
In $\mathbb{R}^{d}$, for each $\varepsilon\leq r$, a ball of radius $r$ has a $\varepsilon$-net with at least $(r/\varepsilon)^{d}$ and at most $(3r/\varepsilon)^{d}$ elements.
@@
Using scaling and translation, it suffices to show the result for the unit ball centered at the origin, $B(\bm{o},1)$, and for $\varepsilon\leq 1$.
@@
$|N|\leq\frac{\operatorname{vol}(B(\bm{o},1+\varepsilon/2))}{\operatorname{vol}(B(\bm{o},\varepsilon/2))}=\frac{V_{d}\cdot\big{(}(1+\varepsilon/2)^{d}\big{)}}{V_{d}\cdot(\varepsilon/2)^{d}}=\left(1+\frac{2}{\varepsilon}\right)^{d}\leq\left(\frac{3}{\varepsilon}\right)^{d},$
where in the last inequality we used that $\varepsilon\leq 1$.
@@ 3 Observations about the problem setup
For each axis, we move $\delta=\lambda(\bm{o})\cdot\varepsilon/(2d)$ along the axis, query the distance to the target, and go back to the origin. The target has to lie in the intersection of the $d$ spheres centered at the points where we queried and also at the sphere centered at the origin. The intersection of these $d+1$ spheres is always a single point, which must be $\bm{t}$. The uniqueness of the point can be seen for example using the lifting to the paraboloid in $\mathbb{R}^{d+1}$ defined by the function $(x_{1},\dots,x_{d})\in\mathbb{R}^{d}\mapsto\left(x_{1},\dots,x_{d},\sum_{i=1}^{d}(x_{i})^{2}\right)\in\mathbb{R}^{d+1}$. Each of those circles corresponds to the intersection of a hyperplane with the paraboloid, and the normals of those $d+1$ planes are linearly independent. Therefore, the intersection of those hyperplanes is a single point, which by construction must lie on the paraboloid.
@@ 4 Upper bound
Assume that we are at point $p_{i}$ and the prediction factor $c^{*}$ is perhaps unknown. Let $c$ be a guess for $c^{*}$. Using a search through a path $\gamma_{i+1}$ of length at most $2(6c)^{d}\cdot\lambda(p_{i})$ we get to one of the following outcomes:
@@
Recall that $|p_{i}\bm{t}|\leq\lambda(p_{i})$. Set $\varepsilon=\tfrac{\lambda(p_{i})}{2c}$ and let $N$ be a $\varepsilon$-net for the ball $B(p_{i},\lambda(p_{i}))$ by Lemma 2.2. Since the target is contained in $B(p_{i},\lambda(p_{i}))$ and $N$ is a $\varepsilon$-net, there is some point $p_{*}\in N$ such that $|p_{*}\bm{t}|\leq\varepsilon=\lambda(p_{i})/2c$. If $c^{*}\leq c$, then we have $\lambda(p_{*})\leq c^{*}\cdot|p_{*}\bm{t}|\leq c^{*}\cdot\lambda(p_{i})/2c\leq\lambda(p_{i})/2$. If $c^{*}>c$, then we have no guaranteed useful bound for $\lambda(p_{*})$.
@@
To bound the length of $\gamma_{i+1}$, we note that any two points in $B(p_{i},\lambda(p_{i}))$ are at distance at most $2\lambda(p_{i})$. We further note that the first edge has length at most $\lambda(p_{i})$ and, if $\gamma_{i+1}$ comes back to $p_{i}$, also the last edge has length at most $\lambda(p_{i})$. Using the bound of Lemma 2.2 we get that the path $\gamma_{i+1}$ has length at most
$\lambda(p_{i})+(|N|-1)\cdot(2\lambda(p_{i}))+\lambda(p_{i})\leq\left(\frac{3\lambda(p_{i})}{\varepsilon}\right)^{d}\cdot 2\lambda(p_{i})=2(6c)^{d}\cdot\lambda(p_{i}).$
@@
Note: It is known [13] that for any set of $n$ points in the $d$-dimensional ball $B(p,r)$ there is a tour of length $r\cdot O\big{(}n^{\frac{d-1}{d}}\big{)}$ visiting them. This implies that we can also use a path $\gamma_{i+1}$ of length $\lambda(p_{i})\cdot O\big{(}|N|^{\frac{d-1}{d}}\big{)}=O((6c)^{d-1})\cdot\lambda(p_{i})$. With this, the dependency on $c$ is slightly better at the expense of having more ugly-looking constants.
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ where the prediction factor $c^{*}>1$ is known. There is a search strategy to reach the target with competitive ratio $2\cdot 6^{d}\cdot(c^{*})^{d+1}$.
@@
For $i=0,1,2\dots$ iteratively, we use Lemma 4.1 with the guessed prediction factor $c=c^{*}$ to obtain a path $\gamma_{i+1}$. As the prediction factor is correct, we always have the outcome in the first item: $\gamma_{i+1}$ finishes at a point $p_{i+1}$ with $\lambda(p_{i+1})\leq\lambda(p_{i})/2$. It follows by induction that for each $i\in\mathbb{N}$ we have $\lambda(p_{i})\leq\lambda(p_{0})/2^{i}$ and therefore $\operatorname{len}(\gamma_{i})\leq 2(6c^{*})^{d}\cdot\lambda(p_{i})\leq 2(6c^{*})^{d}\cdot\lambda(p_{0})/2^{i}$.
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ where the prediction factor $c^{*}$ is unknown. There is a search strategy to reach the target with competitive ratio $(12c^{*})^{d+1}$.
@@
Figure 4: Visualizing the sequence of points $\Big{(}\bigl{(}p^{(j)}_{i}\bigr{)}_{i\in I_{j}}\Bigr{)}_{j\in J}$. When increasing $i$ we move right, when increasing $j$ we move down. Here, we are using $j^{*}=\max(J)$ and $i_{j}=\max(I_{j})$ for all $j\in J$.
@@
Let $j^{*}$ be the largest value that index $j$ takes through the procedure. Note that $j^{*}\leq\lceil\log_{2}c^{*}\rceil$ since $2^{\lceil\log_{2}c^{*}\rceil}$ is an overestimate to $c^{*}$. If we arrive to $j^{*}=\lceil\log_{2}c^{*}\rceil$, from that point on we will always extend the search path using the outcome in the first item of Lemma 4.1. Set $J=\{1,\dots,j^{*}\}$, which is the set of values that index $j$ takes through the procedure.
@@
For each $j\in J$, let $I_{j}$ be the set of indices $i$ such that the point $p^{(j)}_{i}$ is defined and let $i_{j}=\max(I_{j})$. Note that $i_{j^{*}}$ is undefined because $I_{j^{*}}$ is infinite, but $i_{j}$ is defined for all $j<j^{*}$ because $I_{j}$ is finite for all $j<j^{*}$. By construction, the index $i_{j}$ is the first element of $I_{j+1}$, for all $j<j^{*}$. It may happen that, for some $j<j^{*}$, the set $I_{j}$ contains a single element. This happens when $j$ is increased in successive steps and thus $p^{(j)}_{i}=p^{(j+1)}_{i}=p^{(j+2)}_{i}$.
@@
It follows by induction that for all $j\in J$ and all $i\in I_{j}$ we have $\lambda(p^{(j)}_{i})\leq\lambda(p^{(1)}_{0})/2^{i}=\lambda(\bm{o})/2^{i}$. Note that this bound is independent of the index $j$. Therefore
@@
$\forall j\in J,\leavevmode\nobreak\ i\in I_{j}:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{len}(\gamma^{(j)}_{i+1})\leq 2(6\cdot 2^{j})^{d}\cdot\lambda(p^{(j)}_{i})\leq 2(6\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}.$
@@
Let $\gamma$ be the concatenation of the paths $\gamma^{(j)}_{i+1}$, in the same order as we constructed them: first $\gamma^{(1)}_{i+1}$ for increasing $i\in I_{1}$, then $\gamma^{(2)}_{i+1}$ for increasing $i\in I_{2}$, and so on until we reach the infinite sequence $\gamma^{(j^{*})}_{i+1}$ for increasing $i\in I_{j^{*}}$. We then have
@@
$\operatorname{len}(\gamma)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{j=1}^{j^{*}}\sum_{i\in I_{j}}\operatorname{len}(\gamma^{(j)}_{i+1})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{j^{*}}\sum_{i\in I_{j}}2(6\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}.$
@@
Using that for all $j<j^{*}$ the sets $I_{j}$ and $I_{j+1}$ have only $i_{j}=\max(I_{j})$ in common, that $2^{j}\leq c^{*}$ for all $j<j^{*}$, and that $2^{j^{*}}<2c^{*}$, we get
@@
$\operatorname{len}(\gamma)\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \sum_{j=1}^{j^{*}-1}2(6\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i_{j}}}+\sum_{i=0}^{\infty}2(6\cdot 2c^{*})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}$
$\leq\leavevmode\nobreak\ 2\cdot 6^{d}\cdot\lambda(\bm{o})\cdot\sum_{j=1}^{j^{*}-1}(2^{d})^{j}+4(12\cdot c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 2\cdot 6^{d}\cdot\lambda(\bm{o})\cdot\frac{(2^{d})^{j^{*}}}{2^{d}-1}+4(12\cdot c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 4\cdot 3^{d}\cdot\lambda(\bm{o})\cdot(2c^{*})^{d}+4(12\cdot c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 8\cdot 12^{d}\cdot(c^{*})^{d}\cdot\lambda(\bm{o})\leq 12^{d+1}\cdot(c^{*})^{d+1}\cdot|\bm{o}\bm{t}|.$
@@
The path makes an infinite number of straight-line moves. Since for each $i\in\mathbb{N}$, the suffix of the path $\gamma$ after $p^{(j)}_{i+1}$, for $j$ such that $i+1\in I_{j}$, is at distance at most $2\cdot\lambda(p^{(j)}_{i+1})\leq 2\cdot\lambda(\bm{o})/2^{i+1}=\lambda(\bm{o})/2^{i}$ from $\bm{t}$, Lemma 2.1 implies that $\bm{t}$ is the endpoint of the path $\gamma$.

### v2

@@ Abstract
We study the problem of searching for a target at some unknown location in $\mathbb{R}^{d}$ when additional information regarding the position of the target is available in the form of predictions. In our setting, predictions come as approximate distances to the target: for each point $p\in\mathbb{R}^{d}$ that the searcher visits, we obtain a value $\lambda(p)$ such that $|p\bm{t}|\leq\lambda(p)\leq c\cdot|p\bm{t}|$, where $c\geq 1$ is a fixed constant, $\bm{t}$ is the position of the target, and $|p\bm{t}|$ is the Euclidean distance of $p$ to $\bm{t}$. The cost of the search is the length of the path followed by the searcher. Our main positive result is a strategy that achieves $(10c)^{d+1}$-competitive ratio, even when the constant $c$ is unknown. We also give a lower bound of roughly $(c/4)^{d-1}$ on the competitive ratio of any search strategy in $\mathbb{R}^{d}$, assuming that $c\geq 4$.
@@ 1 Introduction
The problem of searching for a target positioned at some unknown location in some region is a classic search game problem that has been well-studied in both fields of Computational Geometry and Operations Research. The problem comes in many different versions, such as linear search (i.e., searching for a target on a line) [9, 6, 7, 16], searching in the plane [8, 21, 17, 4], searching in concurrent rays [15, 16, 4], and searching inside polygonal regions [23, 29]. The book by Alpern and Gal [1] provides an extensive overview of general search games, while Ghosh and Klein [18] survey search problems in planar domains.
@@ 1.1 Problem setup
As we search for the target, we have approximate information about the distance to it from each point that we have visited so far. More precisely, we assume that there is a constant $c\geq 1$ and an unknown function
@@ 1.2 Our contribution
  • We show that for each dimension $d$ and each constant prediction factor $c\geq 1$ there is a search strategy with competitive ratio smaller than $(5c)^{d+1}$. To achieve this, we use $\varepsilon$-nets from metric spaces, also known as $r$-nets, and provide a path of finite length but an infinite number of pieces. This result holds assuming that we know the prediction factor $c$. For unknown prediction factor, a slightly different search strategy leads to a competitive ratio smaller than $(10c)^{d+1}$. These results are given in Section 4.
  • We show that for $c\geq 4$, any deterministic search strategy in $\mathbb{R}^{d}$ with $c$-predictions will have a competitive ratio of at least $(c/4)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$. For this result, we construct an infinite family of $c$-predictions and use a volume argument to give a lower bound on the length of any path that can discern which $c$-prediction from the family is the actual one. The approach is motivated by the techniques used to obtain approximation algorithms for the Euclidean Traveling Salesperson with Neighbourhoods. However, in our setting we have an infinite number of neighbourhoods, one for each $c$-prediction. With a slightly worse constant, the lower bound holds also for randomized search strategies. This lower bound is shown in Section 5.
@@
To our knowledge, this is the first search problem in Euclidean spaces with $d\geq 2$ where a search point reaches a target point with constant competitive ratio; see the discussion below about related work. For our upper bounds we use bounds on the cardinality of $\varepsilon$-nets. For the sake of simplicity in the calculations, we use suboptimal but simple-to-parse estimates in the simplifications and the cardinality of $\varepsilon$-nets. Similarly, in our lower bounds we also use volume estimates that are easier to manipulate. In any case, our upper and lower bounds for the best competitive ratio are still far apart.
@@ 1.3 Related work
In another variant of the problem in the plane by Gal [17], the searcher travels along a path until the target lies on the segment connecting the searcher’s current and starting positions, essentially sweeping around its starting position with an infinitely elastic cord until the target is swept. For this problem, Gal [17] gave a spiral search strategy achieving a competitive ratio of $17.289{\ldots}$, while Langetepe [25] showed that this ratio is optimal.
@@ 2 Notation and preliminaries
The following bound on $\varepsilon$-nets follows from a well-known technique using volumes. We will not use the lower bound, but we include it to show that the upper bound is a reasonable estimate.
@@
In $\mathbb{R}^{d}$, for each $\varepsilon\leq r/2$, a ball of radius $r$ has a $\varepsilon$-net with at least $(r/\varepsilon)^{d}$ and at most $(5r/2\varepsilon)^{d}$ elements.
@@
Using scaling and translation, it suffices to show the result for the unit ball centered at the origin, $B(\bm{o},1)$, and for $\varepsilon\leq 1/2$.
@@
$|N|\leq\frac{\operatorname{vol}(B(\bm{o},1+\varepsilon/2))}{\operatorname{vol}(B(\bm{o},\varepsilon/2))}=\frac{V_{d}\cdot\big{(}(1+\varepsilon/2)^{d}\big{)}}{V_{d}\cdot(\varepsilon/2)^{d}}=\left(1+\frac{2}{\varepsilon}\right)^{d}\leq\left(\frac{5}{2\varepsilon}\right)^{d},$
where in the last inequality we used that $\varepsilon\leq 1/2$.
@@ 3 Observations about the problem setup
For each axis, we move $\delta=\lambda(\bm{o})\cdot\varepsilon/(2d)$ along the axis, query the distance to the target, and go back to the origin. The target has to lie in the intersection of the $d$ spheres centered at the points where we queried and also at the sphere centered at the origin. The intersection of these $d+1$ spheres is always a single point, which must be $\bm{t}$. The uniqueness of the point can be seen for example using the lifting to the paraboloid in $\mathbb{R}^{d+1}$ defined by the function $(x_{1},\dots,x_{d})\in\mathbb{R}^{d}\mapsto\left(x_{1},\dots,x_{d},\sum_{i=1}^{d}(x_{i})^{2}\right)\in\mathbb{R}^{d+1}$. Each of those spheres corresponds to the intersection of a hyperplane with the paraboloid, and the normals of those $d+1$ hyperplanes are linearly independent. Therefore, the intersection of those hyperplanes is a single point, which by construction must lie on the paraboloid.
@@ 4 Upper bound
Assume that we are at point $p_{i}$ and the prediction factor $c^{*}$ is perhaps unknown. Let $c\geq 1$ be a guess for $c^{*}$. Using a search through a path $\gamma_{i+1}$ of length at most $2(5c)^{d}\cdot\lambda(p_{i})$ we get to one of the following outcomes:
@@
Recall that $|p_{i}\bm{t}|\leq\lambda(p_{i})$. Set $\varepsilon=\tfrac{\lambda(p_{i})}{2c}\leq\tfrac{\lambda(p_{i})}{2}$ and let $N$ be a $\varepsilon$-net for the ball $B(p_{i},\lambda(p_{i}))$ given by Lemma 2.2. See Figure 3. Note that
$|N|\leq\left(\frac{5\lambda(p_{i})}{2\varepsilon}\right)^{d}\leq(5c)^{d}.$
Since the target is contained in $B(p_{i},\lambda(p_{i}))$ and $N$ is a $\varepsilon$-net, there is some point $p_{*}\in N$ such that $|p_{*}\bm{t}|\leq\varepsilon=\lambda(p_{i})/2c$. If $c^{*}\leq c$, then we have $\lambda(p_{*})\leq c^{*}\cdot|p_{*}\bm{t}|\leq c^{*}\cdot\lambda(p_{i})/2c\leq\lambda(p_{i})/2$. If $c^{*}>c$, then we have no guaranteed useful bound for $\lambda(p_{*})$.
@@
To bound the length of $\gamma_{i+1}$, we note that any two points in $B(p_{i},\lambda(p_{i}))$ are at distance at most $2\lambda(p_{i})$. We further note that the first edge has length at most $\lambda(p_{i})$ and, if $\gamma_{i+1}$ comes back to $p_{i}$, also the last edge has length at most $\lambda(p_{i})$. Using the bound of Lemma 2.2 for $|N|$, we get that the path $\gamma_{i+1}$ has length at most
$\lambda(p_{i})+(|N|-1)\cdot(2\lambda(p_{i}))+\lambda(p_{i})\leq 2\cdot|N|\cdot\lambda(p_{i})=2(5c)^{d}\cdot\lambda(p_{i}).$
@@
Note: It is known [14] that for any set of $n$ points in the $d$-dimensional ball $B(p,r)$ there is a tour of length $r\cdot O\big{(}n^{\frac{d-1}{d}}\big{)}$ visiting them. This implies that we can also use a path $\gamma_{i+1}$ of length $\lambda(p_{i})\cdot O\big{(}|N|^{\frac{d-1}{d}}\big{)}=O((5c)^{d-1})\cdot\lambda(p_{i})$. With this, the dependency on $c$ is slightly better at the expense of having more ugly-looking constants hidden in the $O$-notation.
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ where the prediction factor $c^{*}>1$ is known. There is a search strategy to reach the target with competitive ratio $4\cdot 5^{d}\cdot(c^{*})^{d+1}$.
@@
For $i=0,1,2\dots$ iteratively, we use Lemma 4.1 with the guessed prediction factor $c=c^{*}$ to obtain a path $\gamma_{i+1}$. As the prediction factor is correct, we always have the outcome in the first item: $\gamma_{i+1}$ finishes at a point $p_{i+1}$ with $\lambda(p_{i+1})\leq\lambda(p_{i})/2$. It follows by induction that for each $i\in\mathbb{N}\cup\{0\}$ we have $\lambda(p_{i})\leq\lambda(p_{0})/2^{i}$ and therefore
[…]
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ where the prediction factor $c^{*}$ is unknown. There is a search strategy to reach the target with competitive ratio $6\cdot 10^{d}\cdot(c^{*})^{d+1}$.
@@
Figure 4: Visualizing the sequence of points $\Big{(}\bigl{(}p^{(j)}_{i}\bigr{)}_{i\in I_{j}}\Bigr{)}_{j\in J}$. When increasing $i$ we move right, when increasing $j$ we move down. Here, we are using $k=\max(J)$ and $i_{j}=\max(I_{j})$ for all $j\in J$.
@@
Let $k$ be the largest value that index $j$ takes through the procedure. Note that $k\leq\lceil\log_{2}c^{*}\rceil$ since $2^{\lceil\log_{2}c^{*}\rceil}$ is an overestimate to $c^{*}$. If we arrive to $k=\lceil\log_{2}c^{*}\rceil$, from that point on we will always extend the search path using the outcome in the first item of Lemma 4.1. Set $J=\{1,\dots,k\}$, which is the set of values that index $j$ takes through the procedure.
@@
For each $j\in J$, let $I_{j}$ be the set of indices $i$ such that the point $p^{(j)}_{i}$ is defined and let $i_{j}=\max(I_{j})$. Note that $i_{k}$ is undefined because $I_{k}$ is infinite, but $i_{j}$ is defined for all $j<k$ because $I_{j}$ is finite for all $j<k$. By construction, the index $i_{j}$ is the first element of $I_{j+1}$, for all $j<k$. It may happen that, for some $j<k$, the set $I_{j}$ contains a single element. This happens when $j$ is increased in successive steps and thus $p^{(j)}_{i}=p^{(j+1)}_{i}=p^{(j+2)}_{i}$.
@@
It follows by induction that, for all $j\in J$ and all $i\in I_{j}$, we have $\lambda(p^{(j)}_{i})\leq\lambda(p^{(1)}_{0})/2^{i}=\lambda(\bm{o})/2^{i}$. Note that this bound is independent of the index $j$. Therefore
@@
$\forall j\in J,\leavevmode\nobreak\ i\in I_{j}:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{len}(\gamma^{(j)}_{i+1})\leq 2(5\cdot 2^{j})^{d}\cdot\lambda(p^{(j)}_{i})\leq 2(5\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}.$
@@
Let $\gamma$ be the concatenation of the paths $\gamma^{(j)}_{i+1}$, in the same order as we constructed them: first $\gamma^{(1)}_{i+1}$ for increasing $i\in I_{1}$, then $\gamma^{(2)}_{i+1}$ for increasing $i\in I_{2}$, and so on until we reach the infinite sequence $\gamma^{(k)}_{i+1}$ for increasing $i\in I_{k}$. We then have
@@
$\operatorname{len}(\gamma)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{j=1}^{k}\sum_{i\in I_{j}}\operatorname{len}(\gamma^{(j)}_{i+1})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{k}\sum_{i\in I_{j}}\left(2(5\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}\right).$
@@
Using that for all $j<k$ the sets $I_{j}$ and $I_{j+1}$ have only $i_{j}=\max(I_{j})$ in common, that $2^{j}\leq c^{*}$ for all $j<k$, that $2^{k}<2c^{*}$, and that $\lambda(\bm{o})\leq c^{*}\cdot|\bm{o}\bm{t}|$, we get
@@
$\operatorname{len}(\gamma)\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \sum_{j=1}^{k-1}\left(2(5\cdot 2^{j})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i_{j}}}\right)+\sum_{i=0}^{\infty}\left(2(5\cdot 2c^{*})^{d}\cdot\frac{\lambda(\bm{o})}{2^{i}}\right)$
$\leq\leavevmode\nobreak\ 2\cdot 5^{d}\cdot\lambda(\bm{o})\cdot\sum_{j=1}^{k-1}(2^{d})^{j}+4(10c^{*})^{d}\cdot\lambda(\bm{o})$
= $=\leavevmode\nobreak\ 2\cdot 5^{d}\cdot\lambda(\bm{o})\cdot\frac{(2^{d})^{k}-2^{d}}{2^{d}-1}+4(10c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 2\cdot 5^{d}\cdot\lambda(\bm{o})\cdot\frac{(2c^{*})^{d}-2^{d}}{2^{d}-1}+4(10c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 2\cdot 5^{d}\cdot\lambda(\bm{o})\cdot 2(c^{*})^{d}+4(10c^{*})^{d}\cdot\lambda(\bm{o})$
$\leq\leavevmode\nobreak\ 4\cdot 5^{d}\cdot(c^{*})^{d}\cdot\lambda(\bm{o})\cdot(1+2^{d})$
$\leq\leavevmode\nobreak\ 4\cdot 5^{d}\cdot(c^{*})^{d}\cdot\lambda(\bm{o})\cdot(\tfrac{3}{2}\cdot 2^{d})$
$\leq\leavevmode\nobreak\ 6\cdot 10^{d}\cdot(c^{*})^{d+1}\cdot|\bm{o}\bm{t}|.$
@@
The path makes an infinite number of straight-line moves. Since for each $i\in I_{k}$, the suffix of the path $\gamma$ after $p^{(k)}_{i+1}$ is at distance at most $2\cdot\lambda(p^{(k)}_{i+1})\leq 2\cdot\lambda(\bm{o})/2^{i+1}=\lambda(\bm{o})/2^{i}$ from $\bm{t}$, Lemma 2.1 implies that $\bm{t}$ is the endpoint of the path $\gamma$.